OpenStudy (anonymous):

8√25√3 Use rational exponents to write as a single radical expression.

OpenStudy (anonymous):

\[\sqrt[2]{8} \sqrt[3]{5} use rational exponents to write as a single radical expression.

OpenStudy (saifoo.khan):

is there a multiply sign in between?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

i can't believe im stuck getting the answer. lol \[\large \sqrt[2]{8} \times \sqrt[3]{5}\] Like that?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Can we get another similar question? Or can we skip this? :(

OpenStudy (anonymous):

If I skip it, I would get it wrong....it is a multiple it choice, it gives 4 different options to it.

OpenStudy (saifoo.khan):

Oh nice, can you provide?

OpenStudy (anonymous):

\[\sqrt[6]{40} \sqrt[6]{12800}, 1/12800, and \sqrt[6]{102400000}\]

OpenStudy (saifoo.khan):

It's \[\Large \sqrt[6]{12800}\]

OpenStudy (anonymous):

\[7\sqrt[3]{5} - 2\sqrt[3]{5}\]

OpenStudy (saifoo.khan):

\[\Large 5 \sqrt[3]{5}\]

OpenStudy (anonymous):

\[3\sqrt{5x+4+2=8}\]

OpenStudy (saifoo.khan):

That can't be possible.

OpenStudy (anonymous):

thats how its written....

OpenStudy (saifoo.khan):

that square root can't be that long.

OpenStudy (anonymous):

it says solve for x

OpenStudy (anonymous):

the option is the solution is not a real number

OpenStudy (saifoo.khan):

\[3 \sqrt{5x+4} + 2 = 8 \]it must be like that?

OpenStudy (saifoo.khan):

Answer is 0

OpenStudy (anonymous):

\[(\sqrt{6+1)} 7\sqrt{6+1)}\]

OpenStudy (anonymous):

mr. saifoo.khan are u tired?

OpenStudy (saifoo.khan):

Oh, sorry i didn't saw that your question is posted. So what is 7?

OpenStudy (anonymous):

the index of the second one

OpenStudy (saifoo.khan):

Can u draw it>

OpenStudy (anonymous):



OpenStudy (anonymous):

in parenthesis

OpenStudy (saifoo.khan):

And what does the question say?

OpenStudy (anonymous):

Multiply. Type an exact answer using radicals as needed.

OpenStudy (saifoo.khan):

\[\sqrt{6+1}^7 * \sqrt{6+1}\]Like that?

OpenStudy (anonymous):

the seven is the index for the second one

OpenStudy (saifoo.khan):

\[\sqrt{6+1} * \sqrt[7]{6+1}\]Like that?

OpenStudy (anonymous):

yes but there is no sign in the middle

OpenStudy (anonymous):

the seven is not raised

OpenStudy (anonymous):